OpenStudy (anonymous):

Simplify: (3√5+2√3) (3√5-2√3)

OpenStudy (anonymous):

9*5-4*3= 45-12= 33

OpenStudy (anonymous):

Thank you! I knew it couldn't be too hard but for some reason I wasn't getting it.